Answer the following problems




A stock is trading at $80 per share. The stock is expected to have a year-end dividend of $4 per share (D1 $4), which is expected to grow at some constant rate g throughout time. The stock’s required rate of return is 14%. If you are an analyst who believes in efficient markets, what is your forecast of g?

1
Expert's answer
2022-03-04T08:29:41-0500

D1 = Stock at the end of the current "14% =\frac{4 (1 + g)}{80 + g}" period

P0 = Current price of stock

g = growth rate

D1 = D0 (1 + g)

= 4 (1 + g)

14% = 4 (1 + g)/80 + g

0.14 = (4 + 4g + 80g)/80

0.14 x 80 = 4 + 84g

11.2 - 4 = 84g

84g = 7.2

g = 7.2/84

g = 0.086 or 8.6%